OpenStudy (anonymous):

20% of what number is 47? Need help

OpenStudy (anonymous):

here you go i hope i helped x/45 = 20/100 45 is on the bottom because it is next to the word "of." *an easy way to remember* ok. now we cross multiply. 100x=900. now solve. 100x/100 900/100= x=9.

OpenStudy (anonymous):

did i help ????

OpenStudy (anonymous):

I gues

OpenStudy (anonymous):

what ever

OpenStudy (anonymous):

(20/100)x=47 20x=47*100 20x=4700 x=4700/20 x=235

OpenStudy (anonymous):

thanks that helped

OpenStudy (anonymous):

no problem